Find the solution of the inhomogeneous heat equation Uxx = U/ +2,0 < x < 1,t> 0; u(0, t) = 0, u(1,t) = 0,t > 0, u(x,0) = x2 –Detail greatly appreciated thankyou
Answer #1

, Их = 4 + 2 оси , +70) »« – Чxx = 10)+ )=0 , щь,+)=0, + 7 — 0 Чех,) = х2-4. Па 14* - 40 и зим Т. или). хор, от te) се, і че)Now by eqn@, e dopusta? mt mtha (s-t) ucxit) - I Din sin (max) e sin falsas no M T n=1 Now using Eqr ② & ④ in above, UCH) = 8ητα) UlXit) = η33 Υ